# beSnake: A routing algorithm for scalable spin-qubit architectures

### v1

@@ Abstract
As quantum computing devices become larger with respect to the number of qubits, the realization of two-qubit interactions becomes more challenging, necessitating innovative and scalable qubit routing solutions. In this work, we introduce beSnake, a novel algorithm specifically designed to address the intricate qubit routing challenges in scalable spin-qubit architectures. The algorithm utilizes shuttle operations that physically move a qubit to an adjacent, unoccupied quantum dot. Unlike traditional methods in superconducting architectures that solely rely on SWAP operations, beSnake incorporates those shuttle operations to optimize the execution time of the circuits and achieve higher operational fidelity and quicker computation times of the routing task itself. By employing a breadth-first search approach, beSnake effectively manages the restrictions created by diverse topologies and positions of qubits. It also dynamically tackles parallelized routing tasks with multiple qubits that act as obstacles with the option to adjust the level of optimization. Our simulations demonstrate beSnake’s advantage over an existing routing solution for random circuits and real quantum algorithms with up to qubits. It shows on average an up to and improvement in gate overhead and depth overhead, respectively, and up to times faster routing time.
@@ I Introduction
For quantum computers to outperform classical systems, enhancements in both hardware and software are essential. The software’s ability to optimally transform hardware-agnostic quantum circuits such that they comply with all operational constraints of the architecture – a process known as mapping – is at the forefront of quantum compiler development [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. One of the tasks of the mapping process is to account for the circuit’s two-qubit interactions and to deal with constraints imposed by the device architectures, notably the limited qubit connectivity. These restrictions are overcome by carefully inserting SWAP gates and thus moving quantum information around the devices so that the needed two-qubit interactions can be performed. This process is also known as qubit routing, and its effectiveness is paramount to maximize operational fidelity of NISQ devices, which are highly prone to errors and have low decoherence times.
@@
SWAP-based qubit routing algorithms have been extensively developed for superconducting quantum devices [11, 12, 13, 14, 15, 16, 17, 18] as these are the most-developed qubit technology when it comes to the number of qubits and system availability in general. Spin-qubit devices, however, are not as mature, and consequently, there are no specialized routing techniques yet that can take advantage of their unique features. The fundamental differentiating aspect of a routing algorithm targeting spin-qubit architectures is the primary method of communication. Here, the shuttling operation is used as a means of communication within the device instead of a SWAP gate, as used in superconducting technology. Whereas a SWAP gate exchanges the quantum state of two qubits, a shuttle operation results in the physical relocation of a qubit. Note that SWAP gates can be implemented in spin-qubit devices; however, shuttling operations are preferred due to their superior operational fidelity and faster execution time. However, this new opportunity also brings unique routing challenges that have not been encountered before. A primary example is the need to have at least one, and preferably more, empty sites for the qubits to be able to move, as two qubits cannot be allocated to the same place at the same time. Note that shuttle operations can also be used for modular ion-trap devices. However, due to fundamental architectural differences, the proposed routing techniques do not apply to spin-qubit devices [19, 20, 21, 22, 23]. Additionally, routing for spin-qubit devices deviates from the conventional notion of enabling two-qubit interactions, it also extends to other types of functions that benefit from the shuttle operation. One of these functions is a fast and low-error single-qubit gate, namely the shuttle-based Z rotation, as used in this [24] crossbar architecture proposal.
@@
On that note, a shuttle-based SWAP routing algorithm for a spin-qubit device has been introduced in [1]. Although it can perform qubit routing in polynomial time in terms of number of qubits, gates, or two-qubit gate percentage, the algorithm is tightly coupled to the strict architectural constraints of a crossbar architecture [24], making it practically unusable for other topologies and constraints. To this date, no routing algorithm has been proposed that fully takes advantage of the shuttle operation, adapts to any architecture, and freely moves qubits around. In this paper, we close this gap by presenting beSnake, a scalable qubit routing algorithm for spin-qubit architectures. In addition to the mentioned characteristics, it is also able to simultaneously handle multiple two-qubit gates and shuttle-based Z rotations, thus aiming at a low depth overhead. beSnake’s core design prioritizes execution speed, thus making it suitable for high qubit counts. In addition, it can be configured on demand regarding the optimization of gate overhead, coming at the expense of total routing time. Other design considerations give beSnake the ability to choose less noisy shuttling paths and, optionally, to intelligently use SWAP operations as long as the architecture supports them. We have extensively evaluated beSnake in several configurations and isolated its features on random quantum circuits. Finally, comparisons with the shuttle-based SWAP algorithm on random and real quantum circuits show on average up to and improvement in gate overhead and depth overhead, respectively, and up to times faster routing time.
@@
The main contribution of this paper is beSnake, the first routing algorithm for scalable spin-qubit architectures. It features the following novel characteristics:
@@
Adapting to various architecture topologies while using noise-aware and adjusting the level of optimization.
@@
Significantly improving over the state-of-the-art (shuttle-based SWAP algorithm) with a considerable decrease in gate and depth overhead for up to qubits.
The remainder of this paper is structured as follows: Section II presents the problem statement regarding routing for spin-qubit architectures with particular use of the shuttle operation. In Section III, we discuss related work and their limitations that include qubit routing techniques used in other quantum technologies as well as algorithms used in classical computing. We present beSnake in Section IV, explain its internal functionality with three representative examples, and discuss special routing cases in detail. Then, in Section V, we thoroughly analyze the performance of beSnake by considering several configurations related to different levels of shortest path optimizations, the SWAP replacement functionality, the behavior on more connected topologies, and various qubit densities. In this section we also compare beSnake with the shuttle-based SWAP routing algorithm on random and real algorithms with up to qubits and discuss the obtained improvements. Finally, we discuss future directions and conclude this paper in Section VI.
@@ II Problem statement
Spin-qubit realizations come with unique physical characteristics that make them a promising technology to scale up quantum computing systems [25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. The fundamental building block offering such characteristics is the quantum dot, in which a confined electron or a hole can define a physical qubit [35]. The spin-qubit can then be controlled electromagnetically through several carefully fabricated gate electrodes around it. These electrodes can enable single-qubit or two-qubit operations by precise pulse sequences across various multi-quantum dot arrangements. Such systems have been extensively studied in 1D and, recently, in 2D array formats [27, 36]. Taking, for instance, the crossbar architecture of [24] shown in Figure 0(a), we can see the different operational lines and sites where spin qubits can be initialized. In this case, they are initialized in a checkerboard pattern. Two-qubit gates are performed between two vertically or horizontally adjacent qubits, meaning that this architecture only supports nearest-neighbor interactions. Therefore, its topology can be represented by a 2D grid as shown in Figure 0(b). Each node (circle) represents a site of the grid, and each edge connecting the nodes represents the possibility of interaction between the two sites. This type of abstract view can be called topology, coupling graph, or layout of the quantum processor.
@@
As the number of quantum dots in a device increases, operating them becomes more complex. As previously mentioned, a key challenge is the requirement for qubits to be adjacent to interact, especially in ever-increasing topology sizes. The problem of bringing the qubits to neighboring positions to execute a two-qubit gate is known as qubit routing. Such a task results in a gate/circuit depth overhead, which is crucial to minimize in NISQ devices as they suffer from high noise rates.
@@
The central distinguishing aspect between spin-qubit and other qubit technologies is the primary communication used for the routing process. As mentioned, in superconducting devices, qubits are ’moved’ employing SWAP gates, in which the quantum state of two neighboring qubits is exchanged. Furthermore, in many superconducting processors, the absence of a native SWAP operation necessitates its decomposition in multiple CNOT gates for its implementation, a process that incurs significant costs. In contrast, for spin-qubit architectures, although SWAP operations are supported as well, shuttling is the preferable communication operation as it offers substantial advantages, including higher operational fidelity and quicker execution time. Moreover, certain quantum dots are deliberately left unoccupied to create free space for facilitating qubit movements, as the shuttle process entails the physical relocation of a qubit to an adjacent vacant quantum dot.
@@
However, the shuttle operations introduce new challenges, especially in the context of a quantum compiler routing process, as highlighted in [1]. One limitation is that at a given moment, there can only be one qubit per site; therefore, it is not physically possible to shuttle a qubit in an already occupied quantum dot. Consequently, there is a need to develop new qubit routing strategies that avoid such conflicts and efficiently move obstacle qubits. Other limitations are related to the unique constraints imposed by the classical control electronics, especially in those with shared control schemes [1], which might create conflicts in certain cases when trying to apply shuttle operations.
@@
The main objective of a routing algorithm is to determine the most efficient path(s) to satisfy a list of, either parallelized or not, two-qubit gates while respecting the architecture’s constraints and avoiding conflicts. This process was shown to be in complexity class NP-complete [37, 38] and thus necessitates a fast and efficient solution, as the goal is to realize scalable architectures that will support thousands of qubits in all sorts of topologies. To date, no existing algorithmic solution has been proposed to tackle this specific path-finding problem in a practically efficient way.
@@ III Related work and its limitations
You can find problems in classical computer science that are similar to qubit routing in spin-qubit architectures. One example is the sliding tile puzzle (STP) [39] that aims to achieve a specified arrangement with the least possible moves within a squared grid by orthogonally sliding tiles into an empty slot. One established method for addressing this problem has been the Interactive Deepening A* (IDA*) algorithm assisted with a precomputed pattern database to increase the accuracy of the admissible heuristic [40, 41]. In more recent work, a 24-tile instance was solved using the Levin Tree Search with Context Models (LTS-CM) [42]. This machine-learning algorithm learned from solved instances of the puzzle and surpasses previous approaches with fewer number of steps. Although these techniques have been very effective, they are not a good candidate for the qubit routing problem described before as: (a) these solutions rely on fine-tuned datasets of particular instances and arrangements of the puzzle. In our problem, creating such a dataset will require large computation resources as we are not aiming for one instance or one specific tile (qubit) configuration but rather a list of goals (i.e., making pairs of qubits adjacent). Another factor to consider is that with more moving directions per location (higher degree), the tree of possible moves expands exponentially in size, as well [43]. However, the presented STP results usually assume these datasets or training and do not disclose the total computational costs. (b) Our problem requires an algorithm that prioritizes speed before optimality (regarding the number of routing steps) to quickly satisfy multiple goals for more than 24 moving parts. (d) As said before, during circuit execution, there can be many steps with multiple two-qubit gates.
@@
Another potentially fitting formulation is the multi-agent path-finding (MAPF), which has been studied extensively for multi-robot systems [44]. Each agent navigates to a target destination through various techniques while avoiding collisions with other agents or obstacles. Agents are sequentially executing a plan consisting of a specific goal set of destinations while trying to minimize the traveled distance during a goal. We have identified, however, the following key differences to our problem: (a) As shown in [44], classical MAPF approaches become slow and unsuccessful for large numbers of agents, especially with obstacles. (b) Our problem is not limited to a single set of goals (i.e., a set of two-qubit gates) but a plan with multiple goals in each step. Therefore, ultimately, the focus is minimizing the cost function for the entire plan and each goal. (c) In our problem, qubits can be viewed as agents or dynamic obstacles. In addition, only a subset of agents (qubits) need to complete goals at each step of a plan, and these goals are not target locations but conditions that could be satisfied in many locations. Currently, there is a lack of efficient strategies specifically tailored to address this particular problem, especially when dealing with a large number of agents, as highlighted in [45, 46, 47].
@@
In the quantum world, several qubit routing techniques have been proposed, but mostly for superconducting and trapped ion platforms as they are the most advanced in terms of qubit counts. Routing for spin qubits resembles the operation of a quantum charge-coupled ion trap device (QCCD) [48], another promising scalable architecture. In QCCD, trap regions are dedicated to tasks such as temporary storage and processing. Communication between ions of different regions is then implemented with shuttles through a shared channel where collision avoidance optimization techniques are used [48, 19]. However, these techniques fundamentally differ as they are made specifically for the QCCD’s unique regions and topology. The particular structure of a QCCD device results in different constraints for moving qubits or parallelizing operations compared to 2D grid spin qubit lattices. Hence, their routing strategies cannot be implemented in spin qubit platforms even though both use shuttling [20, 22].
@@
In superconducting systems, routing involves consecutive state exchanges by means of SWAP gates along a chosen path until the qubit pair is adjacent. The most commonly used methods for qubit routing are based on heuristic search algorithms or reinforcement learning techniques [19, 20, 21, 22, 23]. The only requirement needed to find this path is the presence of neighboring locations within the topology where the two qubits can interact. In spin-qubit architectures, routing entails the physical relocation of qubits instead of state swapping; thus, all qubit positions play an important role during routing, and there needs to be at least one empty location to be able to move qubits around. When routing in superconducting devices, the predominant approach is to identify the shortest paths between qubit operands, concurrently minimizing costs across several parameters. These parameters include circuit gate and depth overhead, as well as noise-aware path selection [19, 20, 21, 22, 23]. These can be optimization goals for a spin-qubit routing algorithm as well, but finding the shortest path is not enough because other qubits might block the way. The problem can become even more complex when considering routing for multiple two-qubit gates at the same time while trying to avoid conflicts and satisfy operational constraints of the architecture [1].
@@
Shuttle-based SWAP (SBS) was conceptualized as the first routing algorithm for large-scale spin qubit architectures [1]. In particular, it was tailored to the unique constraints of the crossbar architecture [24], which necessitated the maintenance of the checkerboard pattern of physical qubits to achieve a fast compilation process. One of the main drawbacks of the complex architectural crossbar constraints was the parallelization limitations imposed on shuttle operations. As a consequence, SBS can only, yet efficiently, route one two-qubit gate at a time on grid topologies for that particular crossbar architecture. As extensively discussed in [1], this severely impacts the compilation overhead for high qubit counts and high two-qubit gate percentages. In this paper, we are extending the work on spin qubit routing to fully utilize and explore the benefits of shuttling-based movement of qubits, assuming a more generalized architecture with less constraints. We propose beSnake, a qubit routing algorithm for large-scale spin qubit processors, which can handle any combination and size of parallelized Z and two-qubit gates in any topology, as opposed to SBS. Furthermore, beSnake is designed to be more flexible and nearly matches the speed of SBS, which opens opportunities for more complex routing tasks. We have highlighted their main differences in Table 1. Based on the early stage of spin-qubit technologies, beSnake is introduced as a concrete baseline routing algorithm for future scalable architectures that use shuttling.

### v2

@@ Abstract
As quantum computing devices increase in size with respect to the number of qubits, two-qubit interactions become more challenging, necessitating innovative and scalable qubit routing solutions. In this work, we introduce beSnake, a novel algorithm specifically designed to address the intricate qubit routing challenges in scalable spin-qubit architectures. Unlike traditional methods in superconducting architectures that solely rely on SWAP operations, beSnake also incorporates the shuttle operation to optimize the execution time and fidelity of quantum circuits and achieves fast computation times of the routing task itself. Employing a simple breadth-first search approach, beSnake effectively manages the restrictions created by diverse topologies and qubit positions acting as obstacles, for up to 72% qubit density. It also has the option to adjust the level of optimization and to dynamically tackle parallelized routing tasks, all the while maintaining noise awareness. Our simulations demonstrate beSnake’s advantage over an existing routing solution on random circuits and real quantum algorithms with up to qubits, showing an average improvement of up to in gate overhead and in depth overhead, and up to times faster routing times.
@@ I Introduction
The quantum software’s ability to optimally transform hardware-agnostic quantum circuits such that they comply with all operational constraints of the architecture – a process known as mapping – is at the forefront of quantum compiler development right now [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. One of the tasks of the mapping process is to account for the circuit’s two-qubit interactions and to deal with constraints imposed by the device architectures, notably the limited qubit connectivity. These restrictions are overcome by carefully inserting SWAP gates and thus moving quantum information around the devices – usually in nearest-neighbor positions – so that the needed two-qubit gates can be performed. This particular process is also known as qubit routing, and its effectiveness is paramount to maximize the operational fidelity of Noisy Intermediate-Scale Quantum (NISQ) [11] devices, which are highly prone to different kinds of errors.
@@
SWAP-based qubit routing algorithms have been extensively developed for superconducting quantum devices [12, 13, 14, 15, 16, 17, 18, 19] as these are the most-developed qubit technology when it comes to the number of qubits and system availability in general. Spin-qubit devices, however, are not as mature, and consequently, there are no specialized routing techniques yet that can take advantage of their unique features. The fundamental differentiating aspect of a routing algorithm targeting spin-qubit architectures is the primary method of communication. Here, the shuttling operation is used as a means of communication within the device instead of SWAP gates used in superconducting technology. Whereas a SWAP gate exchanges the quantum state of two qubits, a shuttle operation results in the physical relocation of a qubit. Note that SWAP gates can also be implemented in spin-qubit devices; however, shuttling operations are preferred due to their superior operational fidelity and faster execution time.
Having said that, the shuttle operation brings unique routing challenges that have not been encountered before. A primary example is the need to have at least one, and preferably more, empty sites for the qubits to be able to move, as two qubits cannot be allocated to the same place at the same time. Of course, this is not new for modular ion-trap devices, as they are also using shuttles to move ions between traps. However, due to their fundamental architectural differences with spin-qubit devices, the proposed routing techniques are not applicable [20, 21, 22, 23, 24]. Additionally, routing for spin-qubit devices deviates from the conventional notion of only enabling two-qubit interactions. It also extends to other types of operations benefiting from the shuttle operation, which necessitates innovative solutions able to handle the different cases. One of these operations is a fast and low-error single-qubit shuttle-based Z rotation, as used in this [25] crossbar architecture proposal.
@@
On that note, a shuttle-based SWAP routing algorithm for a spin-qubit device has been introduced in [1]. Although it can perform qubit routing in polynomial time in terms of number of qubits, gates, or two-qubit gate percentage, the algorithm is tightly coupled to the strict architectural constraints of a crossbar architecture [25], making it practically unusable for other architectures. To this date, no routing algorithm has been proposed that fully takes advantage of the shuttle operation, adapts to any architecture, and freely moves qubits around the topology. In this paper, we close this gap by presenting beSnake, a novel routing algorithm for scalable spin-qubit architectures. In addition to the mentioned characteristics, it is also able to simultaneously handle multiple two-qubit gates and shuttle-based Z rotations, thus aiming at a low circuit depth overhead. beSnake’s core design prioritizes execution speed, thus making it suitable for high qubit counts. In order to further reduce the computation time, it can be configured accordingly at the expense of slightly higher gate overhead. Other design considerations give beSnake the ability to choose less noisy shuttling paths and, optionally, use SWAP operations, assuming that the architecture supports them.
We have extensively evaluated beSnake by routing random generated quantum circuits that stress-test all its capabilities. Finally, comparisons with the Shuttle-based SWAP (SBS) algorithm [1] on random and real quantum circuits show on average an up to and improvement in gate overhead and depth overhead, respectively, and up to times faster routing time.
@@
The main contribution of this paper is beSnake, the truly first routing algorithm for scalable spin-qubit architectures. It features the following novel characteristics:
@@
Adapting to various architecture topologies while being noise-aware and able to adjust the level of optimization.
@@
Significantly improving over the state-of-the-art (the Shuttle-based SWAP routing algorithm) with a considerable decrease in gate and depth overhead for up to qubits.
The remainder of this paper is structured as follows: Section II presents the problem statement regarding routing for spin-qubit architectures with particular use of the shuttle operation. In Section III, we discuss related work and their limitations, including qubit routing techniques used in other quantum technologies as well as algorithms used in classical computing. We present beSnake in Section IV, explain its internal functionality with three representative examples, and discuss special routing cases in detail. Then, in Section V, we thoroughly analyze the performance of beSnake by considering several configurations related to different levels of shortest path optimizations, the SWAP replacement functionality, the behavior on more connected topologies, and on various qubit densities. In the same section, we also compare beSnake with the Shuttle-based SWAP routing algorithm [1] on random and real algorithms with up to qubits and discuss the obtained improvements. Finally, we discuss future directions and draw conclusions in Section VI.
@@ II Problem statement
Spin-qubit realizations come with unique physical characteristics that make them a promising technology to scale up quantum computing systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The fundamental building block offering such characteristics is the quantum dot, in which a confined electron or a hole can define a physical qubit [36]. The spin-qubit can then be controlled electromagnetically through several carefully fabricated gate electrodes around it. These electrodes can enable single-qubit or two-qubit operations by precise pulse sequences across various multi-quantum dot arrangements. Such systems have been extensively studied in 1D and, recently, in 2D array formats [28, 37]. Taking, for instance, the crossbar architecture of [25] shown in Fig. 1(a), we can see the different operational lines and sites where spin qubits can be initialized. In this case, they are initialized in a checkerboard pattern. Two-qubit gates are only performed between two vertically or horizontally adjacent qubits, meaning that this architecture, or any architecture for that matter, only supports nearest-neighbor interactions. Therefore, its topology can be represented by a 2D grid, as shown in Fig. 1(b). Each node (circle) represents a site of the grid, and each edge connecting the nodes represents the possibility of interaction between the two sites. This type of abstract view is interchangeably called either topology, coupling graph, or layout of the quantum processor.
@@
The problem of bringing the qubits to neighboring positions to execute a two-qubit gate is known as qubit routing. During the NISQ era, it is crucial for the routing process to strive for gate and depth overhead minimization of quantum circuits due to devices suffering from high noise rates. As the number of quantum dots in spin-qubit processors increases in the future, operating them will be even more complex. Qubit routing, then, becomes a key challenge, especially in ever-increasing topology sizes.
@@
Turning now our attention to other qubit technologies, the central distinguishing aspect of spin-qubit quantum processors is the primary communication method used for the routing process. As mentioned, in superconducting devices, qubits are “moved” employing SWAP gates, in which the quantum state of two neighboring qubits is exchanged. Furthermore, in many superconducting processors, the absence of a native SWAP operation necessitates decomposition in multiple CNOT gates for its implementation, a process that can incur significant errors. In contrast, in spin-qubit architectures, although SWAP operations are supported as well, shuttling is the preferable communication operation as it offers substantial advantages, including higher operational fidelity and quicker execution time. Moreover, certain quantum dots are deliberately left unoccupied to create free space for facilitating qubit movements, as the shuttle process entails the physical relocation of a qubit to an adjacent vacant quantum dot.
@@
However, this introduces new challenges, especially in the context of a quantum compiler routing process, as highlighted in [1]. This is because, at any given moment, there can be only one qubit per site; therefore, it is not physically possible to shuttle a qubit in an already occupied quantum dot. Consequently, there is a need to develop new qubit routing strategies that avoid such conflicts and efficiently move obstacle qubits away. Other limitations are related to the unique constraints imposed by the classical control electronics, especially in those with shared control schemes [1], which might create conflicts in certain cases when trying to apply parallelized shuttle operations.
@@
Broadly speaking, the main objective of a routing algorithm is to determine the most efficient path(s) to satisfy a list of, either parallelized or not, two-qubit gates while respecting the architecture’s constraints and avoiding conflicts. Starting with a defined topology, initial physical and virtual-to-physical qubit placement, the algorithm receives a sequence of two-qubit gates, which, in most cases, cannot be executed directly. Therefore, the primary objective of a router is to identify and correctly integrate qubit movement operations (i.e., shuttles or SWAPs) into the existing gate sequence to ensure that all two-qubit gates can be executed. The overarching goal is to minimize the additional circuit overhead required to achieve this. In the end, the final revised sequence, which includes these movement operations, constitutes the output of our routing algorithm. This process was shown to be in complexity class NP-complete [38, 39] and thus necessitates a fast solution, as the goal is to use it for scalable architectures that will support thousands of qubits in all sorts of topologies. To date, no algorithmic solution has been proposed to tackle this specific path-finding problem in a practically efficient way for the plethora of unique characteristics of spin-qubit architectures.
@@ III Related work and its limitations
One can find problems in classical computer science that are similar to qubit routing in spin-qubit architectures. One example is the sliding tile puzzle (STP) [40] that aims to achieve a specified arrangement with the least possible moves within a squared grid by orthogonally sliding tiles into an empty slot. One established method for addressing this problem has been the Interactive Deepening A* (IDA*) algorithm assisted with a precomputed pattern database to increase the accuracy of the admissible heuristic [41, 42]. In more recent work, a 24-tile STP instance was solved using the Levin Tree Search with Context Models (LTS-CM) [43]. This machine-learning algorithm learned from solved instances of the puzzle and surpasses previous approaches with fewer number of steps. Although these two techniques and others have been very effective, they are not a good candidate for the qubit routing problem for scalable spin-qubit devices: (a) these solutions heavily rely on fine-tuned datasets of particular instances and arrangements of the puzzle. In our problem, creating such a dataset will require large computation resources as we are not aiming for one instance or one specific tile (qubit) configuration but rather a list of goals (i.e., making pairs of qubits adjacent). (b) Another factor to consider with point (a) is that with more moving directions per location (higher node degree), the tree of possible moves expands exponentially in size [44], as well. The presented STP results in the literature usually assume these datasets or the training phase and do not disclose the total computational costs. However, upon taking a closer look, it becomes obvious that as the number of qubits increases, these approaches become practically intractable. Our problem requires an algorithm that prioritizes speed before optimality (regarding the number of routing steps) to quickly satisfy multiple goals for more than 24 moving parts. (c) As previously mentioned, during circuit executions there are many steps, some of which involve parallelized two-qubit gates. This contrasts with the objective of aiming for a single tile pattern in the STP problem.
@@
Another potentially fitting formulation is the multi-agent path-finding (MAPF) problem, which has been studied extensively for multi-robot systems [45]. Each agent, representing an independent robotic system, navigates to a target destination through various techniques while avoiding collisions with other agents or obstacles. Agents are sequentially executing a plan consisting of specific goal destinations while trying to minimize the traveled distance within a goal. We have identified, however, the following key differences to our problem: (a) As shown in [45], classical MAPF approaches become slow and unsuccessful for large numbers of agents, especially with the existence of obstacles. (b) Our problem is not limited to a single set of goals (i.e., a set of two-qubit gates) but a plan with multiple goals in each step. Therefore, ultimately, the focus is on minimizing some cost function for the entire plan (representing a quantum circuit), as it is crucial to incur the least errors possible. (c) In our problem, qubits can be viewed as agents and as dynamic obstacles at the same time. In addition, only a subset of agents (qubits) need to complete goals at each step of the plan, and these goals are not target locations but conditions that could be satisfied in multiple locations. Currently, there is a lack of strategies specifically tailored to address this particular problem in a short amount of time, especially when dealing with a large number of agents, as highlighted in [46, 47, 48].
@@
In the quantum world, several qubit routing techniques have been proposed, but mostly for superconducting and trapped ion platforms as they are the most advanced in terms of qubit counts. Routing for spin qubits resembles a similar functionality of quantum charge-coupled ion trap devices (QCCD) [49], another promising scalable architecture. In QCCD, trap regions are dedicated to specific tasks such as temporary storage or processing. Communication between ions of different regions is thus implemented with shuttles through a shared channel where collision avoidance optimization techniques are used [49, 20]. However, these techniques fundamentally differ because they are made specifically for the QCCD’s unique regions and topology. The unique structure of a QCCD device imposes a distinct set of constraints on moving qubits and parallelizing operations, different from to those found in 2D spin qubit topologies. Hence, these routing strategies cannot be implemented in spin qubit platforms even though both use shuttling [21, 23].
@@
In superconducting systems, routing involves consecutive state exchanges by means of SWAP gates along a chosen path until the desired qubit pair is adjacent. The most commonly used methods for qubit routing in such devices are based on heuristic search algorithms or reinforcement learning techniques [20, 21, 22, 23, 24]], although recently more theoretical proposals have been introduced for bounding the minimal number of SWAPs needed [50, 51]. The only requirement to determine a viable path is the presence of at least one pair of neighboring locations within the topology where the two qubits can interact. In spin-qubit architectures, on the other hand, routing entails the physical relocation of qubits instead of state swapping. Therefore, within the notion of bringing two qubits closer to each other until adjacent, other qubit positions play an important role as well.
When routing in superconducting devices, the predominant approach is to identify the shortest paths between qubit operands, concurrently minimizing costs across several parameters. These parameters include circuit gate and depth overhead, as well as noise-aware path selection [20, 21, 22, 23, 24]. These can be optimization goals for a spin-qubit routing algorithm as well, but finding the shortest path is not enough because other qubits might block the way. The problem can become even more complex when considering routing for multiple two-qubit gates at the same time while trying to avoid conflicts and satisfy operational constraints of the architecture [1].
@@
The Shuttle-based SWAP (SBS) algorithm was conceptualized as the first routing algorithm for large-scale spin qubit crossbar architectures [1]. In particular, it was tailored to the unique and complex constraints of this crossbar architecture [25], which necessitated the maintenance of the checkerboard pattern of physical qubits to achieve a fast compilation process. One of the main drawbacks of these constraints was the parallelization limitations imposed on shuttle operations. As a consequence, SBS can only, yet efficiently, route one two-qubit gate at a time on grid topologies for that particular crossbar architecture. As extensively discussed in [1], this severely impacts the mapping overhead for high qubit counts and high two-qubit gate percentages.
In this paper, we are extending the work on spin qubit routing to fully utilize and explore the benefits of shuttling-based movement of qubits, assuming a more generalized architecture with fewer constraints. We propose beSnake, a qubit routing algorithm for large-scale spin qubit processors, which can handle any combination and number of parallelized shuttle-based Z and two-qubit gates in any topology. Furthermore, beSnake is designed to be more flexible, which opens opportunities for more complex routing tasks. We are highlighting their main differences in Table 1. Based on the early stage of spin-qubit technologies, beSnake is introduced as a concrete baseline routing algorithm for future scalable architectures that use shuttling.